1. Deluxe Company expects to pay a dividend of $2 per share at the end of year 1. $3 per share at the end of year 2. and then be sold for $32 per share at the end of year 2. If the required rate of return on the stock is 15%, what is the current value of the stock?

a. $32.17

b. $32

c. $29.18

d. $28.20
Answer #1

According to dividend discount model, the current value of a stock is the sum of the present value of the future dividends and the present value of the terminal value:

Following is the cash flow

Year 1 2
Cashflow 2 3+32 = 35

C1 = 2, C2 = 35

Interest rate = 15%

The present value of the cash flows C1 and C2 is:

PV = C1/(1+r) + C2/(1+r)2 = 2/(1+15%) + 35/(1+15%)2

PV = 2/1.15 + 35/1.152 = 1.73913043478261 + 26.4650283553875 = 28.2041587901701

Answer -> $28.20